# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v1

@@ 3.4 Rendering and Optimization
where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i^{th}$ sample. $L_{o}$ is the entropy on the opacity by [25], defined as:
@@
$L_{r}$ is L1 regularization on the spatial features in the current branch as:
@@
We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 for alleviating foggy effects.
@@ 4.1 Datasets
DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated. DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames. It contains five 300-frame videos and a 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS). To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds. Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS. Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background. Please refer to the supplementary material for more details.
@@ 4.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. Please see the supplementary material for more details.
@@ 4.3 Results and Comparisons
offline DyNeRF [28] 300 29.58 0.020 0.083 7 days 28MB >100GB
NeRF-T [28] 300 28.45 0.023 0.10 - - >100GB
Interp-NN [34] 300 29.88 - 0.096 2 days 20MB >100GB
HexPlane [7] 300 31.57 0.016 0.089 96 min 200MB 62GB
K-Planes [17] 300 31.63 - - 110 min 200MB >100GB
MixVoxels-T [45] 300 30.71 0.024 0.162 15 min 500MB >100GB
4DGS [54] 300 32.01 - 0.055 6 hours 4000MB <10GB
HyperReel [1] 50 31.1 - 0.096 9 hours 360MB 18GB
NeRFplayer [43] 300 30.29 - 0.152 5.5 hours - >100GB
online StreamRF [27] 1 28.26 0.029 0.226 75 min 5000MB <10GB
HexPlane-PR 10 24.03 0.081 0.244 100 min 318MB 14GB
HexPlane-GR 10 25.17 0.050 0.272 125 min 318MB 14GB
INV [48] 1 28.72 - 0.121 40 hours 336MB <10GB
CD-NGP (Ours) 10 30.23 0.029 0.198 75 min 113MB 14GB
4.3.1 Standard-length sequences.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the released output of [27] for DSSIM and LPIPS metrics. [27, 48] are methods trained per-frame. [43] is trained on the whole video sequence but streamable upon inference. 4DGS [54] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
@@
We compare our method with both online and offline baselines. As demonstrated in Tab. 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation [48], $\times4$ smaller than the voxel-based continual representation [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like previous methods [27, 48], our method achieves fast convergence and compact model size simultaneously: it leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 54], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
Figure 3 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based continual learning methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4. As illustrated in Fig. 6, our method reaches a new balance between rendering quality, model size, and training time while remaining memory-friendly, in which we use metrics on the flame salmon scene for fair comparison. The perceptual consistency is measured by $1-$ LPIPS, and the speed is the reciprocal of the training time. The compactness and RAM-Efficiency are mapped by $-log$ function from the model size and memory occupation respectively.
@@
4.3.2 Long sequences.
We also provide quantitative comparisons on our proposed long multi-view video dataset in Tab. 3 and Fig. 6 to further show the advantage in scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. The offline baseline [7] produces unstable results on one of the 5 scenes in our dataset, so we only use the other 4 scenes to obtain the metrics of [7]. As shown in Tab. 3, our method reduces memory and storage consumption by at least 4×4$\times$4 × with comparable quality to the offline baseline. Compared with the results in Tab. 1, the performance gap in storage consumption is even larger in long video sequences, because the auxiliary branches consume the majority of the parameters and the complexity in storage is reduced to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. The PSNR curves in Fig. 6 further demonstrate the stability of our method. The offline baseline produces the best quality. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than at the end. Our method resists the problem by parameter isolation and ensures the quality of the whole sequence. We also provide comparisons between rendered results in Fig. 8. The regularization method in Fig. 7(a) reconstructs the TV screen in the background erroneously. The replay method in Fig. 7(b) fails to recover the details. Ours in Fig. 7(d) produces comparable quality to the offline baseline in Fig. 7(c).
@@
Figure 6: PSNR curves of the seminar scene. The solid line and dotted line denotes offline and online methods respectively.
@@ 4.4 Ablation studies
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Tab. 3. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches. Please note the whole model sizes include the sizes of the occupancy grid and are thus larger than those deduced from parameter numbers. Larger hash tables alleviate hash collisions and lead to better results. However, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 8. Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
Table 2: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset. denotes the method trained offline. Metrics are measured on the first 1200 frames. Compared baselines are trained with a chunk size of 300 frames.
@@
HexPlane-PR 20.90 0.286 1272MB 12GB
HexPlane-GR 24.82 0.282 1272MB 12GB
CD-NGP (ours) 26.50 0.201 201MB 12GB
HexPlane [7] 28.32 0.117 800MB 74GB
@@
Table 5: Ablation study on regularization weight $\lambda_{r}$. $10^{-3}$ is the default value.
[…]
$10^{-5}$ 30.20 0.0269 0.198
$10^{-4}$ 30.18 0.0268 0.196
$10^{-3}$ 30.23 0.0269 0.198
$10^{-2}$ 30.09 0.0279 0.197
1 29.66 0.0313 0.205
The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones. We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Tab. 4. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 6 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation.
Table 4 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 7 and 6 with vector concatenation and denote them as Voxel-C and MERF-C in Tab. 4. The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Tab. 3, the Voxel-C representation in Tab. 4 almost performs as well as the bulky $(18,18)$ configuration in Tab. 3 and reduces the model size by $80\%$.
We also provide ablation studies on the regularization weights of auxiliary spatial encodings. We apply regularizations on $\Psi_{k}(x)$ with different $\lambda_{r}$ and show the results in Tab. 5. The regularization is of slight improvement, and our method is robust within a wide range of $\lambda_{r}$. Please refer to the supplementary material for more ablation studies.
5 Conclusion and Limitation
Appendix 0.A Detailed Results on DyNeRF Dataset
We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Tab. 6.
@@ Appendix 0.A Detailed Results on DyNeRF Dataset
Appendix 0.B More Implementation Details
0.B.1 Parameter numbers and streaming efficiency.
We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the proposed CD-NGP and MERF representations within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Tab. 4 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Tab. 4. We use $2^{P-2}$-length hash tables for the Plane representation in Tab. 4 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Tab. 4 are set as $(6,4)$ for better results.
The detailed parameter numbers and bandwidth consumption are thus listed in Tab. 7. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
Table 7: The parameter numbers of each module of the 30-branch representations in Tab. 4. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).

### v3

@@ 2.4 Rendering and optimization
where $s_{i}$ denotes the normalized ray distance along the ray, $w_{i}=T_{i}\alpha_{i}$ is the weight of the $i$-th sample.
$L_{o}$ in Eq. 8 is the entropy on the opacity by [25], defined as:
@@
$L_{r}$ in Eq. 8 is the L1 regularization on the spatial features in the current branch as:
@@
We set $\lambda_{r}$ as 0.001 to encourage leverage of the base hash encoder. Based on [25], we set a larger $\lambda_{d},\lambda_{o}$ as 0.005 to alleviate foggy effects.
2.5 Continual learning pipeline.
Algorithm 1 Continual learning pipeline
0:  The whole representation for dynamic scenes.
[…]
2:  while $N_{chunk}$ do
[…]
5:     if $k\leavevmode\nobreak\ \mathbf{mod}\leavevmode\nobreak\ T_{episode}\leavevmode\nobreak\ =0$  then
[…]
7:     else
[…]
10:     end if
11:     Train the $k$-th model branch $(\Psi_{k},\Gamma_{k},\theta_{1}^{(k)},\theta_{2}^{(k)})$ for $\eta$ steps.
[…]
13:  end while
@@ 3.1 Datasets
DyNeRF dataset [28] is the most prevailing dataset for dynamic view synthesis that provides high-quality multi-camera videos, on which most of the current state-of-the-art methods are evaluated. DyNeRF dataset contains 6 individual cooking scenes with varied illumination, food topological changes, and transient effects like flames. It contains five 300-frame videos and one 1200-frame video, with up to 21 synchronized camera views at the resolution of $2704\times 2028$ at 30 frames per second (FPS).
Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution. The lengths of the scenes are set as 300 frames and the refresh rates are set as 30 FPS.
To further prove the viability of our method on long video sequences, we provide a new dataset containing 5 dynamic scenes at $3840\times 2160$ resolution, the lengths of which are at least 200 seconds. Similar to the DyNeRF dataset [28], scenes in our dataset are recorded by 20 GoPro cameras at 30 FPS. We synchronize the videos by environmental sound, and we put chessboard images in the scene to help calibrate the cameras and estimate the camera poses. Compared with those in the DyNeRF dataset, the dynamics are much harder, including large motions of human bodies and transient effects in the background. We collect the dataset for evaluating the scalability of our method and further research. We have checked the videos to make sure they do not contain controversial or offensive content before using them for research purposes. The faces in the videos may contain identifiable information, but the information will neither be used for any other purposes except for research nor be disclosed to any third party. All videos are collected with the explicit consent of all the participants. The view in the upper center is selected for validation.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following [7], we exclude the unsynchronized coffee martini scene and conduct experiments on the others. Following most of the state-of-the-art methods [28, 27, 17, 45, 43] we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We use $1280\times 720$ resolution for our dataset to fit into memory. We use the importance sampling scheme in [28] at a batch size of 1024 rays. The parameters are optimized by Adam [23], with $\epsilon=10^{-15},\beta=(0.9,0.96)$. The learning rate is set as 0.001 and scheduled by cosine function following [25]. We set the size of the hash table in the first branch as $2^{19}$ and the size of the hash table in the subsequent branches as $2^{14}$, and we set the structural parameters $(L,F)$ in the spatial hash encoders as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for 18k iterations, and for each of the subsequent branches we train for 3k iterations using the previous branch for initialization. We use LeakyReLU activation in the MLP regressors. For long video sequences, we assign 30 chunks as an episiode. The MLP regressors are re-trained from initialization in the first branch of each episode to adapt to recent changes. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision.
We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene, and we use $(64,2048)$ for the plane representation.
Following MERF [38] for static scenes, we use a smaller hash table for the MERF and the MERF-C than pure voxel hash grid [31] representation in Table 5 to ensure the compactness and scalability of our method. Denoting the lengths of hash tables along pure voxels in a branch of CD-NGP as $2^{P}$, the lengths of hash tables in the MERF and the MERF-C are set as $2^{P-3}$ and $2^{P-4}$ for voxel and plane encoders respectively. The number of the parameters of the MERF representation is thus reduced to $1-2^{-3}-3\times 2^{-4}=31.25\%$ of the voxel hash grids. Since the hash tables take up the majority of the parameters, the model size is reduced from $21M$ to $7.5M$ as in Table 5. We use $2^{P-2}$-length hash tables for the Plane representation in Table 5 to balance the quality and the model size. The structural configurations $(L,F)$ of the plane representation in Table 5 are set as $(6,4)$ for better results.
@@ 3.3 Results and comparisons
DyNeRF [28] $\checkmark$ 300 29.58 0.020 0.083 7 days 28MB >100GB
NeRF-T [28] $\checkmark$ 300 28.45 0.023 0.10 - - >100GB
Interp-NN [34] $\checkmark$ 300 29.88 - 0.096 2 days 20MB >100GB
HexPlane [7] $\checkmark$ 300 31.57 0.016 0.089 96 min 200MB 62GB
K-Planes [17] $\checkmark$ 300 31.63 - - 110 min 200MB >100GB
MixVoxels-T [45] $\checkmark$ 300 30.71 0.024 0.162 15 min 500MB >100GB
4DGS [55] $\checkmark$ 300 32.01 - 0.055 6 hours 4000MB <10GB
HyperReel [1] $\checkmark$ 50 31.1 - 0.096 9 hours 360MB 18GB
NeRFplayer [43] $\checkmark$ 300 30.29 - 0.152 5.5 hours - >100GB
StreamRF [27] $\checkmark$ 1 28.26 0.029 0.226 75 min 5000MB <10GB
HexPlane-PR $\checkmark$ 10 24.03 0.081 0.244 100 min 318MB 14GB
HexPlane-GR $\checkmark$ 10 25.17 0.050 0.272 125 min 318MB 14GB
INV [48] $\checkmark$ 1 28.72 - 0.121 40 hours 336MB <10GB
CD-NGP (Ours) $\checkmark$ 10 30.23 0.029 0.198 75 min 113MB 14GB
3.3.1 DyNeRF dataset.
Table 1 shows the quantitative comparisons between CD-NGP and other state-of-the-art methods on DyNeRF dataset. We use the officially released metrics. The methods in [28, 34] are trained on 8 V100 GPUs or equally 4 A100 GPUs. The methods in [28, 34, 43] have no official implementations available. We are unable to obtain the missing quality metrics in [17] because of their huge memory occupation in the default resolution. HexPlane [7] uses resolution $(1024,768)$ in their implementation, so the memory occupation is much smaller than other offline methods. We use the released output of [27] for DSSIM and LPIPS metrics. StreamRF [27] and INV [48] are trained per-frame. NeRFplayer [43] is trained offline but streamable upon inference. 4DGS [55] caches the frames onto disk for training to alleviate memory occupation with a trade-off on training speed. The reconstruction quality is measured by PSNR, DSSIM, and LPIPS [57] (by AlexNet [24]) following common practices [28, 7, 45]. To prove the viability of our parameter-isolation-based method, we apply the previous state-of-the-art method HexPlane [7] to continual learning setting with parameter-regularization (HexPlane-PR) and generative replay (HexPlane-GR) as additional baselines.
@@
We compare our method with both online and offline baselines. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently-configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also achieves $40\%$ less model size ($0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches with comparable quality.
@@
Figure 3 provides qualitative comparisons. Compared with offline methods in Fig. 3(a), Fig. 3(b) and the online method in Fig. 3(c), our model is much smaller without large quality loss. Compared with regularization-based and replay-based continual learning methods in Fig. 3(d) and Fig. 3(e) respectively, our method achieves significantly higher quality. Results of other scenes including depth maps are shown in Fig. 4.
(a) GT
(b) HexPlane
(c) MixVoxels
(d) INV
(e) StreamRF
(f) Ours
Figure 5: Comparison of reconstruction quality on the trimming scene in the Meetroom dataset. Our method produces comparable quality to both offline and online state-of-the-art methods but with much less memory and storage cost.
3.3.2 Meetroom dataset.
Table 2 and Fig. 5 shows the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6 is the PSNR curve of the compared methods on the trimming scene. As shown in Table 2, HexPlane-GR and HexPlane-PR fail and produce unsatisfactory results on the meetroom dataset, we thus exclude them from Fig. 6. Our method outperforms the implicit MLP-based online method [48] in terms of quality, speed, and model size significantly. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed, and consumes much less memory as in Fig. 5. We use the released configuration of [27] for the meetroom dataset. The full model of [27] results in much worse metrics than those reported in their paper, so we only report the metrics of their pilot model.
@@
Figure 6: PSNR curves of the compared methods on the trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively.
Table 2: Results on the Meetroom dataset.
[…]
Mixvoxels $\checkmark$ 26.96 0.027 0.1034 15min 500MB
HexPlane $\checkmark$ 24.71 0.043 0.2206 100min 200MB
HexPlane-GR $\checkmark$ 19.05 0.099 0.3999 100min 318MB
HexPlane-PR $\checkmark$ 18.63 0.112 0.3589 100min 318MB
StreamRF $\checkmark$ 26.48 0.029 0.1699 75min 1710MB
INV $\checkmark$ 23.87 0.041 0.1195 40hours 336MB
Ours $\checkmark$ 26.01 0.039 0.1906 77min 113MB
3.3.3 Long sequence dataset and scalability analysis.
We also provide quantitative comparisons on our proposed long multi-view video dataset in Table 3 and Fig. 7 to further show the advantage in scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 3, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the whole model shares one base branch and the following auxiliary branches occupy the majority of the parameters. We thus set the hash table size in the auxiliary branches to $2^{12}LF$ and leverage the re-use of the features from the base branch to reduce the model size to $\mathcal{O}(N_{chunk}\cdot 2^{12}LF)$. .
The PSNR curves in Fig. 7 further demonstrate the stability of our method. The compared online methods suffer from catastrophic forgetting: the PSNRs at the beginning of the chunks are much lower than that in the end. The problem is properly eliminated by our method through parameter isolation, and our method continuously produce high-quality results for the whole sequence. We also provide comparisons between the rendered results in Fig. 8. The regularization method in Fig. 8(a) reconstructs the TV screen in the background erroneously. The replay method in Fig. 8(b) fails to recover the details. Ours in Fig. 8(f) produces comparable quality to the offline baseline in Fig. 8(c).
Figure 7: PSNR curves of the seminar scene in our long video dataset. The solid lines and dotted lines denote offline and online methods respectively.
@@ 3.4 Ablation studies
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the parameter numbers of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 4. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the other branches. Since the whole model contains hash encoders, MLP regressors, and the occupancy grids. The sizes of the models are thus larger than the simple combinations of hash table and MLP regressors. We use the default configuration $(19,14)$ in the following experiments. Although larger hash tables alleviate hash collisions, the configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 9. Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The advantage comes from our asymmetric hash table designs: the base branch with a large hash table recovers more environment patterns, and the auxiliary branches could thus use fewer parameters to cope with recent changes only.
Table 3: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset. denotes the method trained offline. Metrics are measured on the first 1200 frames. The compared HexPlane baselines are trained with a chunk size of 300 frames.
@@
HexPlane-PR 20.90 0.286 1272MB 8 hours 12GB
HexPlane-GR 24.82 0.282 1272MB 8 hours 12GB
HexPlane 28.32 0.117 800MB 8 hours 74GB
StreamRF 24.99 0.266 13GB 4 hours 12GB
INV 27.07 0.086 1344MB 160 hours 12GB
Ours 27.15 0.190 201MB 10 hours 12GB
@@
3.4.2 Spatial representations and scalability analysis.
The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones. To prove the flexibility and scalability of our method, we provide ablation studies on different spatial representations. We adopt the Eq. 5 to the MERF representation [38], defined as:
$\Psi(\mathbf{x})=\bigoplus_{m=1}^{L}\leavevmode\nobreak\ P_{m}(x,y,z)+Q_{mxy}(x,y)+Q_{myz}(y,z)+Q_{mzx}(z,x)$ (12)
where $Q_{mij}(i,j)$ denotes the bi-linearly interpolated value from the 4 nearby vertices in the orthogonally projected planes w.r.t. axes $(i,j)$ at the $m$-th resolution. Removing the $P_{m}(x,y,z)$ term leads to a pure plane representation like [17, 21].
We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Table 5. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5. By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
Table 5 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 6 and 12 with vector concatenation and denote them as Voxel-C and MERF-C in Table 5. The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Table 4, the Voxel-C representation in Table 5 almost performs as well as the bulky $(18,18)$ configuration in Table 4 and reduces the model size by $80\%$. We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 6.
@@
The detailed parameter numbers and bandwidth consumption are thus listed in Table 7. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
Table 7: The parameter numbers of each module of the 30-branch representations in Table 5. $\mathbf{B_{min}}$ and $\mathbf{B_{avg}}$ are minimum and average bandwidth costs in MB/frame. ##\## denotes the number of parameters per branch (10 frames).